# Any-stepsize Gradient Descent for Separable Data under Fenchel–Young Losses

### v1

@@ Appendix B Proof of Theorem 5
where the invertibility of $g$ can be verified through the differentiability of $g$ as in Lemma 14 (by relying upon $\phi^{\prime\prime}>0$ in Assumption 3A), and the last identity follows by the definition of the convex conjugate. Plugging this into $C_{\phi}$ defined in Lemma 14, we see that it matches $C_{\phi}$ in (3). Thus, we have proven all statements. ∎
Appendix C Proof of Theorem 7
@@
Most of the results in this section have already been provided in Wu et al. [55]. We restate the statements and proofs here to make the paper self-contained, and moreover, simplify the statements from the NTK setup to the linear-model case to highlight the essential structures.
@@
Suppose Assumption 2 and consider a convex and nonincreasing loss $\ell$ satisfying Assumptions 3B and 3C. For every $t\geq 1$,
@@
Proof.
By invoking Lemma 11 with the choice of $\theta$
@@
$\leq\frac{\rho(\gamma^{2}\eta t)}{\gamma^{2}\eta t}+\frac{1}{2\eta t}\left(\theta+\frac{\eta C_{g}}{2\gamma}\right)^{2},$ (Lemma 8)
@@
Suppose Assumption 2 and consider a convex and nonincreasing loss $\ell$ satisfying Assumptions 3B and 3C. Then, we have
@@
$\geq\left\langle{\mathbf{w}_{t}},{\mathbf{w}_{*}}\right\rangle-\frac{\gamma\eta}{n}\sum_{i=1}^{n}\ell^{\prime}(\left\langle{\mathbf{w}_{t}},{\mathbf{z}_{i}}\right\rangle)$ (Assumption 2 and note $-\ell^{\prime}(\cdot)\geq 0$)
@@
where the last inequality is due to the parameter bound in Lemma 15. ∎
@@
Consider a loss satisfying 4A. Suppose there exists $s<T$ such that
@@
Proof.
We first show that Claim 1 implies Claim 2. By Assumption 4A, we have
@@
Next, we prove Claim 1 by induction. The base case $t=s$ holds by Assumption 4A as follows:
@@
Consider a nonincreasing loss $\ell$ satisfying Assumptions 3B and 4A. Suppose there exists $s<T$ such that
@@
Proof.
By the lemma assumption, we can apply Lemma 17 for $s$ onwards. Therefore, we have for $k\geq 0$,
@@
$\leq\|\mathbf{w}_{s+k}-\mathbf{u}\|^{2}+2\eta\left\langle{\nabla L(\mathbf{w}_{s+k})},{\mathbf{u}-\mathbf{w}_{s+k}}\right\rangle+\frac{4}{3}\eta L(\mathbf{w}_{s+k}).$ (by (11))
@@
The assumption $L(\mathbf{w}_{s})\leq\ell(0)/n$ allows us to apply Lemma 18, so $\left\langle{\mathbf{w}_{s}},{\mathbf{z}_{i}}\right\rangle\geq 0$ and thus
@@
Telescoping the sum from 00 to $t-1$ and rearranging, we get
@@
$\leq 3\frac{\rho(\gamma^{2}\eta t)}{\gamma^{2}\eta t}+\frac{3\|\mathbf{w}_{s}-\mathbf{u}\|^{2}}{2\eta t}.$ (Lemma 8)
@@
By Lemma 17, $L(\mathbf{w}_{t})$ is nonincreasing for $t\geq s$, and thus we have
@@
Suppose Assumption 2 and consider a convex and nonincreasing loss $\ell$ that satisfies Assumptions 3B and 3C. Define
@@
Proof.
Applying Lemma 15 with $t=\tau$, we have
@@
Suppose Assumption 2 and consider a nonincreasing loss $\ell$ satisfying Assumptions 3B and 3C, and 4B. Furthermore, assume $\eta\geq 1$. Then, there exists $C>0$ depending on $C_{e}$, $C_{g}$, $C_{\beta}$, $\ell(0)$, and $n$ such that the following holds. Let
@@
Proof.
Under Assumption 4B, we have
@@
Applying Lemma 16, we have
@@
where the last inequality is due to Assumption 4B. The above implies $\left\langle{\mathbf{w}_{s}},{\mathbf{z}_{i}}\right\rangle\geq 0$ since $g(\cdot)$ is nonincreasing. Thus, we can apply Assumption 4B for $\left\langle{\mathbf{w}_{s}},{\mathbf{z}_{i}}\right\rangle$ and get
@@
Appendix D Separation margin and self-bounding property

### v2

@@ Appendix B Proof of Theorem 5
where the invertibility of $g$ can be verified through the differentiability of $g$ as in Lemma˜13 (by relying upon $\phi^{\prime\prime}>0$ in Assumption˜3A), and the last identity follows by the definition of the convex conjugate. Plugging this into $C_{\phi}$ defined in Lemma˜13, we see that it matches $C_{\phi}$ in (3). Thus, we have proven all statements. ∎
Appendix C Extension to the stochastic gradient descent
We discuss the extension of Theorem˜5 to the stochastic setup. We consider the constant-stepsize online stochastic gradient descent (SGD) as follows:
$\mathbf{w}_{t+1}\coloneqq\mathbf{w}_{t}-\eta\nabla L_{t}(\mathbf{w}_{t}),\quad\text{where}\quad L_{t}(\mathbf{w})\coloneqq\ell(\left\langle{\mathbf{w}},{y_{t}\mathbf{x}_{t}}\right\rangle),\quad t\geq 0,$ (SGD)
for a loss function $\ell\colon\mathbb{R}\to\mathbb{R}_{\geq 0}$. Here, $(\mathbf{x}_{t},y_{t})_{t\geq 0}$ are independently and identically distributed according to the following assumption.
Assumption 4.
Assume the training data $(\mathbf{x}_{t},y_{t})_{t\geq 0}$ are independent copies of $(\mathbf{x},y)$ following a distribution such that
  1. 1.
    $\|\mathbf{x}\|\leq 1$, and $y_{i}\in\left\{{\pm 1}\right\}$, almost surely;
  2. 2.
    there is $\gamma>0$ and a unit vector $\mathbf{w}_{*}$ such that $\left\langle{\mathbf{w}_{*}},{\mathbf{z}_{t}}\right\rangle\geq\gamma$ for $\mathbf{z}_{t}\coloneqq y_{t}\mathbf{x}_{t}$, almost surely.
Proposition 14.
Suppose Assumption˜4 and consider (SGD) with stepsize $\eta>0$ and $\mathbf{w}_{0}=\mathbf{0}$ under a Fenchel–Young loss $\ell$ satisfying Assumption˜3, and additionally having separation margin $m>0$. For arbitrary $\delta,\bar{\varepsilon}\in(0,1)$, define $(\alpha,C_{\phi})$ as in Eq.˜3, for which we assume $\alpha,C_{\phi}\in(0,\infty)$. In addition, for arbitrary $\varepsilon\in(0,\bar{\varepsilon})$, define
$M_{\eta,\gamma}\coloneqq\ell\left(-\frac{4m+\eta C_{g}}{\gamma}\right)\text{~~~and~~~}t^{\circ}\coloneqq\max\left\{{\frac{32M_{\eta,\gamma}^{2}\ln(1/\delta)}{\varepsilon^{2}},\frac{8M_{\eta,\gamma}\ln(1/\delta)}{\varepsilon}}\right\}.$
Then, if we run (SGD) with $T$ iterations such that
$T>N\cdot t^{\circ},\text{~~~where~~~}N\coloneqq\frac{2^{\alpha}}{C_{\phi}\gamma^{2}}\left(\frac{4m}{\eta}+C_{g}\right)\varepsilon^{-\alpha},$
then we have $\min_{t\in[T]}\operatorname*{\mathbb{E}}[L_{t}(\mathbf{w}_{t})]\leq\varepsilon$ with probability at least $(1-\delta)^{N}$.
Before proving Proposition˜14, several auxiliary lemmas are presented. Since the proof of Proposition˜14 closely follows Theorem˜5, the following Lemmas˜15 and 16 are almost identical to the deterministic versions (Lemmas˜11 and 19, respectively), and hence we omit the proofs.
Lemma 15.
Suppose Assumption˜4 and consider (SGD) with any stepsize $\eta>0$ under a Fenchel–Young loss that satisfies Assumption˜3A. Moreover, assume that there exists $k\in[0,t-1]$ such that we have a non-trivial lower bound $g(\left\langle{\mathbf{w}_{k}},{\mathbf{z}_{k}}\right\rangle)\geq g_{\min}>0$. Then, we have
$\gamma\eta g_{\min}\leq\left\langle{\mathbf{w}_{t}},{\mathbf{w}_{*}}\right\rangle-\left\langle{\mathbf{w}_{0}},{\mathbf{w}_{*}}\right\rangle.$
Lemma 16.
Suppose Assumption˜4 and consider (SGD) with any stepsize $\eta>0$ under a convex and nondecreasing loss $\ell$ satisfying Assumptions˜3B and 3C. For every $t\geq 1$, we have
The following concentration result is an additional argument that we need in the stochastic case.
Lemma 17.
Consider (SGD) with any stepsize $\eta>0$ under a Fenchel–Young loss that satisfies Assumptions˜3A and 3C. Let $Z_{0},\dots,Z_{t-1}$ be the independent copies of data $\mathbf{z}=y\mathbf{x}$ following Assumption˜4. We introduce the filtration $\left\{{\mathcal{F}_{t}}\right\}_{t\geq 0}$, where $\mathcal{F}_{k}$ is a $\sigma$-algebra defined on $Z_{1},\dots,Z_{k}$, and let $H_{k}\coloneqq\ell(\left\langle{\mathbf{w}_{k}},{Z_{k}}\right\rangle)$ be a random variable, where $\mathbf{w}_{k}$ is an $\mathcal{F}_{k-1}$-measurable random variable. We write $S_{t}\coloneqq\sum_{k=0}^{t-1}H_{k}$ for a random variable standing for the accumulated loss, and let $\overline{S}_{t}\coloneqq S_{t}/t$. Moreover, we introduce the following assumptions:
  • (Bounded mean) $\mu_{k}\coloneqq\operatorname*{\mathbb{E}}[H_{k}|\mathcal{F}_{k-1}]\leq M$ for $k=0,\dots,t-1$.
  • (Bounded variance) $\mathrm{Var}(H_{k}|\mathcal{F}_{k-1})\leq\sigma^{2}$ for $k=0,\dots,t-1$.
Then, for any $\varepsilon>0$ and $\delta\in(0,1)$, if we have
$\frac{\operatorname*{\mathbb{E}}[S_{t}]}{t}>\varepsilon\text{~~~and~~~}t\geq\max\left\{{\frac{32\sigma^{2}\ln(1/\delta)}{\varepsilon^{2}},\frac{8M\ln(1/\delta)}{\varepsilon}}\right\},$ (10)
then we have $\overline{S}_{t}>\varepsilon/2$ with probability at least $1-\delta$.
Proof..
First, we apply a type of the martingale inequality, Freedman’s inequality [22, Theorem 1.6]. Since $\operatorname*{\mathbb{E}}[H_{k}-\mu_{k}|\mathcal{F}_{k-1}]$ is the mean-zero martingale difference with the bounded variance $\sigma^{2}$, Freedman’s inequality is applicable. Then, we have
$\Pr\left\{{S_{t}\leq\operatorname*{\mathbb{E}}[S_{t}]-u}\right\}\leq\exp\left[-\frac{u^{2}}{2(Mu+t\sigma^{2})}\right].$
Equivalently, we have the following inequality with probability at least $1-\delta$:
$S_{t}$ $>\operatorname*{\mathbb{E}}[S_{t}]-\sqrt{M^{2}\ln^{2}(1/\delta)+2t\sigma^{2}\ln(1/\delta)}-M\ln(1/\delta)$
$>\operatorname*{\mathbb{E}}[S_{t}]-\sqrt{2t\sigma^{2}\ln(1/\delta)}-2M\ln(1/\delta).$
Dividing by $t$, we have
$\overline{S}_{t}>\frac{\operatorname*{\mathbb{E}}[S_{t}]}{t}-\sqrt{\frac{2\sigma^{2}\ln(1/\delta)}{t}}-\frac{2M\ln(1/\delta)}{t}>\varepsilon-\frac{\varepsilon}{4}-\frac{\varepsilon}{4}=\frac{\varepsilon}{2},$
where we used the conditions (10) at the second inequality. Thus, the desired inequality is shown. ∎
Now we are ready to prove Proposition˜14. Overall, the proof consists of the perceptron argument and the concentration property.
Proof of Proposition˜14.
The first step is to establish the concentration property. To apply Lemma˜17, we confirm the bounded moment conditions. For the mean $\operatorname*{\mathbb{E}}[H_{k}|\mathcal{F}_{k-1}]=\operatorname*{\mathbb{E}}[\ell(\left\langle{\mathbf{w}_{k}},{Z_{k}}\right\rangle)|\mathcal{F}_{k-1}]$, where $\mathcal{F}_{k-1}$ is the $\sigma$-algebra defined on $\left\{{Z_{l}}\right\}_{l=1}^{k-1}$, we have
$\operatorname*{\mathbb{E}}[\ell(\left\langle{\mathbf{w}_{k}},{Z_{k}}\right\rangle)|\mathcal{F}_{k-1}]$ $\leq\operatorname*{\mathbb{E}}[\ell(-\|\mathbf{w}_{k}\|)|\mathcal{F}_{k-1}]$ ($\ell$ is nonincreasing and $\|Z_{k}\|\leq 1$)
$\leq\ell\left(-\frac{4m+\eta C_{g}}{\gamma}\right)$ (Lemmas˜15 and 9)
$\eqqcolon M_{\eta,\gamma}.$
For the variance $\mathrm{Var}(H_{k}|\mathcal{F}_{k-1})$, we similarly have
$\mathrm{Var}(H_{k}|\mathcal{F}_{k-1})$ $=\operatorname*{\mathbb{E}}[H_{k}^{2}|\mathcal{F}_{k-1}]-\operatorname*{\mathbb{E}}[H_{k}|\mathcal{F}_{k-1}]^{2}$
$\leq\operatorname*{\mathbb{E}}[H_{k}^{2}|\mathcal{F}_{k-1}]$
$\leq\ell\left(-\frac{4m+\eta C_{g}}{\gamma}\right)^{2}$
$=M_{\eta,\gamma}^{2}.$
Note that these moment bounds hold uniformly for any $k$. By plugging $M=M_{\eta,\gamma}$ and $\sigma^{2}=M_{\eta,\gamma}^{2}$ into Lemma˜17, if we have
$\frac{1}{t^{\circ}}\sum_{k=t_{0}}^{t_{0}+t^{\circ}-1}\operatorname*{\mathbb{E}}\left[L_{k}(\mathbf{w}_{k})\right]>\varepsilon\text{~~~and~~~}t^{\circ}\geq\max\left\{{\frac{32M_{\eta,\gamma}^{2}\ln(1/\delta)}{\varepsilon^{2}},\frac{8M_{\eta,\gamma}\ln(1/\delta)}{\varepsilon}}\right\},$ (11)
then we have
$\frac{1}{t^{\circ}}\sum_{k=t_{0}}^{t_{0}+t^{\circ}-1}L_{k}(\mathbf{w}_{k})>\frac{\varepsilon}{2}\text{~~~with probability at least $1-\delta$.}$
Let us use this concentration argument. Split the interval $[0,T]$ into length-$t^{\circ}$ sub-intervals (for $t^{\circ}$ satisfying (11)) such that
$0,T]=\underbrace{[0,t^{\circ}-1]}_{\eqqcolon\mathcal{I}_{1}}\sqcup\underbrace{[t^{\circ},2t^{\circ}-1]}_{\eqqcolon\mathcal{I}_{2}}\sqcup\underbrace{[2t^{\circ},3t^{\circ}-1]}_{\eqqcolon\mathcal{I}_{3}}\dots\sqcup\underbrace{[(N-1)t^{\circ},Nt^{\circ}-1]}_{\eqqcolon\mathcal{I}_{N}}\sqcup[Nt^{\circ},T].$ (12)
Consider the scenario where $\operatorname*{\mathbb{E}}[L_{k}(\mathbf{w}_{k})]>\varepsilon$ holds for all $k\in[0,T]$, and focus on an arbitrary sub-interval $\mathcal{I}_{l}$. Since we have $\frac{1}{t^{\circ}}\sum_{k\in\mathcal{I}_{l}}\operatorname*{\mathbb{E}}[L_{k}(\mathbf{w}_{k})]>\varepsilon$, the concentration argument implies that $\frac{1}{t^{\circ}}\sum_{k\in\mathcal{I}_{l}}L_{k}(\mathbf{w}_{k})>\varepsilon/2$ with probability at least $1-\delta$. This further indicates the high-probability existence of $k_{l}\in\mathcal{I}_{l}$ such that $L_{k_{l}}(\mathbf{w}_{k_{l}})>\varepsilon/2$. In this case, we have $g(\left\langle{\mathbf{w}_{k_{l}}},{\mathbf{z}_{k_{l}}}\right\rangle)\geq g(\ell^{-1}(\varepsilon/2))$ for this specific $k_{l}\in\mathcal{I}_{l}$, which can be seen in the same manner as the proof of Theorem˜5. Since this concentration argument does not depend on the sub-interval choice $\mathcal{I}_{l}$, there exists a set of indices $\left\{{k_{l}}\right\}_{l\in[N]}$ such that each of $g(\left\langle{\mathbf{w}_{k_{l}}},{\mathbf{z}_{k_{l}}}\right\rangle)\geq g(\ell^{-1}(\varepsilon/2))$ holds with probability at least $1-\delta$.
Next, we invoke the perceptron argument. By combining Lemmas˜15 and 16 with the choice $g_{\min}=g(\ell^{-1}(\varepsilon/2))$, we have
$\gamma\eta g\left(\ell^{-1}\left(\frac{\varepsilon}{2}\right)\right)\cdot N\leq\left\langle{\mathbf{w}_{T}},{\mathbf{w}_{*}}\right\rangle\leq\|\mathbf{w}_{T}\|\leq\frac{4\sqrt{\rho(\gamma^{2}\eta T)}+\eta C_{g}}{\gamma}\leq\frac{4m+\eta C_{g}}{\gamma},$
with probability at least $(1-\delta)^{N}$, where we additionally used Lemma˜9 at the last inequality. By applying Lemma˜13 at the left-most side, we have
$\gamma\eta C_{\phi}\cdot\left(\frac{\varepsilon}{2}\right)^{\alpha}\cdot N\leq\frac{4m+\eta C_{g}}{\gamma},$
which implies
$N\leq\frac{2^{\alpha}}{C_{\phi}\gamma^{2}}\left(\frac{4m}{\eta}+C_{g}\right)\varepsilon^{-\alpha}\quad(\eqqcolon N(\varepsilon)).$
Therefore, if $N>N(\varepsilon)$ (or $T>N(\varepsilon)\cdot t^{\circ}$), with probability at least $(1-\delta)^{N}$, we have $\operatorname*{\mathbb{E}}[L_{k}(\mathbf{w}_{k})]\leq\varepsilon$ for some $k\in[0,T]$.
Finally, we need verify that $C_{\phi}$ defined in Lemma˜13 matches (3), but we skip it because it can be confirmed in the same manner as in the proof of Theorem˜5. ∎
C.1 Comparison between GD and SGD
Whereas the iteration complexity for GD given by Corollary˜6 is $T=\Omega(\varepsilon^{-\alpha})$, the iteration complexity for SGD given by Proposition˜14 is $T=\Omega(\varepsilon^{-(\alpha+2)})$. Hence, the SGD rate is significantly slower than the GD case. This deterioration is because we can observe a non-trivial lower bound $g(\left\langle{\mathbf{w}_{k}},{\mathbf{z}_{k}}\right\rangle)$ after every $t^{\circ}=\Omega(\varepsilon^{-2})$ steps. As in the GD case, we need $N=\Omega(\varepsilon^{-\alpha})$ such non-trivial lower bounds, and hence the total iteration number amounts to $Nt^{\circ}=\Omega(\varepsilon^{-(\alpha+2)})$. While this apparently looks a big bottleneck, note that the GD rate is $T\gtrsim n\varepsilon^{-\alpha}$ if we explicitly write the $n$-dependency. Since (SGD) consumes only one fresh sample at every update (while (GD) consumes $n$ samples), the extra complexity $N=\Omega(\varepsilon^{-2})$ appearing in the SGD case compensates for this gap of sample sizes. The GD/SGD rates are comparable in this sense.
Appendix D Phase transition of large-stepsize GD
Assumption 5.
  1. A.
  2. B.
Theorem 18 ([63]).
Consider (GD) with stepsize $\eta>0$ and initialization $\mathbf{w}_{0}=\mathbf{0}$ under a loss $\ell$ satisfying Assumptions˜3B and 3C, and 5A. Let $T$ be the maximum number of steps. Then, we have the following:
  • The EoS phase. For every $t>0$, we have
  • The stable phase. If $s<T$ is such that
    If $\tau\leq T$, (13) holds for some $s\leq\tau$. Moreover, if $\ell$ additionally satisfies Assumption˜5B and $\eta\geq 1$, there exists $C_{2}>0$ that depends on $C_{e}$, $C_{g}$, $C_{\beta}$, $\ell(0)$, and $n$ such that $\tau$ is improved as follows:
    $\tau\coloneqq\frac{C_{2}}{\gamma^{2}}\max\left\{{\eta,n}\right\}.$
@@ Appendix D Phase transition of large-stepsize GD
The proof consists of Lemma˜19 (the EoS phase), Lemma˜23 (the stable phase), and Lemmas˜24 and 25 (phase transition time), respectively. Most of the results in this section have already been provided in Wu et al. [63]. We restate the statements and proofs here to make the paper self-contained, and moreover, simplify the statements from the NTK setup to the linear-model case to highlight the essential structures.
@@
Suppose Assumption˜2 and consider a convex and nonincreasing loss $\ell$ satisfying Assumptions˜3B and 3C. For every $t\geq 1$, we have
@@
Proof..
By invoking Lemma˜10 with the choice of $\theta$
@@
$\leq\frac{\rho(\gamma^{2}\eta t)}{\gamma^{2}\eta t}+\frac{1}{2\eta t}\left(\theta+\frac{\eta C_{g}}{2\gamma}\right)^{2},$ (Lemma˜7)
@@
Suppose Assumption˜2 and consider a convex and nonincreasing loss $\ell$ satisfying Assumptions˜3B and 3C. Then, we have
@@
$\geq\left\langle{\mathbf{w}_{t}},{\mathbf{w}_{*}}\right\rangle-\frac{\gamma\eta}{n}\sum_{i=1}^{n}\ell^{\prime}(\left\langle{\mathbf{w}_{t}},{\mathbf{z}_{i}}\right\rangle)$ (Assumption˜2 and note $\ell^{\prime}(\cdot)\geq 0$)
@@
where the last inequality is due to the parameter bound in Lemma˜19. ∎
@@
Consider a loss satisfying 5A. Suppose there exists $s<T$ such that
@@
Proof..
We first show that Claim 1 implies Claim 2. By Assumption˜5A, we have
@@
Next, we prove Claim 1 by induction. The base case $t=s$ holds by Assumption˜5A as follows:
@@
Consider a nonincreasing loss $\ell$ satisfying Assumptions˜3B and 5A. Suppose there exists $s<T$ such that
@@
Proof..
By the lemma assumption, we can apply Lemma˜21 for $s$ onwards. Therefore, we have for $k\geq 0$,
@@
[…]
@@
The assumption $L(\mathbf{w}_{s})\leq\ell(0)/n$ allows us to apply Lemma˜22, so $\left\langle{\mathbf{w}_{s}},{\mathbf{z}_{i}}\right\rangle\geq 0$ and thus
@@
Telescoping the sum from 0 to t1t-1 and rearranging, we get
@@
$\leq 3\frac{\rho(\gamma^{2}\eta t)}{\gamma^{2}\eta t}+\frac{3\|\mathbf{w}_{s}-\mathbf{u}\|^{2}}{2\eta t}.$ (Lemma˜7)
@@
By Lemma˜21, $L(\mathbf{w}_{t})$ is nonincreasing for $t\geq s$, and thus we have
@@
Suppose Assumption˜2 and consider a convex and nonincreasing loss $\ell$ that satisfies Assumptions˜3B and 3C. Define
@@
Proof..
Applying Lemma˜19 with $t=\tau$, we have
@@
Suppose Assumption˜2 and consider a nonincreasing loss $\ell$ satisfying Assumptions˜3B and 3C, and 5B. Furthermore, assume $\eta\geq 1$. Then, there exists $C>0$ depending on $C_{e}$, $C_{g}$, $C_{\beta}$, $\ell(0)$, and $n$ such that the following holds. Let
@@
Proof..
Under Assumption˜5B, we have
@@
Applying Lemma˜20, we have
@@
where the last inequality is due to Assumption˜5B. The above implies $\left\langle{\mathbf{w}_{s}},{\mathbf{z}_{i}}\right\rangle\geq 0$ since $g(\cdot)$ is nonincreasing. Thus, we can apply Assumption˜5B for $\left\langle{\mathbf{w}_{s}},{\mathbf{z}_{i}}\right\rangle$ and get
@@
Appendix E Separation margin and self-bounding property